OpenStudy (anonymous):

The problem is "The mean of a set of 15 numbers is 9. One of the numbers is 10. What is the mean of the other 14 numbers?" Textbook gives an answer of 125 but I don't get it

OpenStudy (anonymous):

x stands for total of fifteen numbers added together x/15 = 9 --->(15) x/15 = (15)(9)...x = 135 The total of the 15 numbers was 135. Now if you take away the 10 you will have 14 numbers. (x - 10)/14 = mean Now substitute 135 in for x... (135 - 10)/14 = 125/14 = 8.9 (rounded = 9) Maybe I did this wrong, but it seems that 125 is the total of the 14 numbers not the mean.